# Natural super-orbitals representation of many-body operators

### v1

@@ Abstract
We introduce the concept of natural super-orbitals for many-body operators, defined as the eigenvectors of the one-body super-density matrix associated with a vectorized operator. We first analyze the general analytical properties of these objects in various contexts, including the time-evolution operator of non-interacting Hamiltonians and Haar-random unitaries. We then perform a numerical investigation of the natural super-orbitals corresponding to both the time-evolution operator and a time-evolved local operator, focusing on two many-body systems: the fermionic $t\text{-}V$ chain and an impurity model, using tensor network simulations. Our results reveal that the $t\text{-}V$ model lacks a preferred super-orbital basis, while in the impurity model, the occupations of the natural orbitals for both operators decay exponentially at all times. This indicates that only a small number of orbitals contribute significantly to quantum correlations, enabling a compact matrix-product-operator representation in the natural orbital basis. Finally, we examine the spatial spread of the natural orbitals for time-evolved local operators in the impurity model and show that the complexity of this operator in the natural orbital basis saturates over time. This new framework opens the door to future research that leverages the compressed structure of operators in their natural super-orbital basis, enabling for instance the computation of out-of-time-order correlators in large interacting systems over extended time scales.
@@ I Introduction
The paper is organized as follows: In Section II, we review the properties of the one-body density matrix for quantum states, focusing on the features we aim to generalize to operators. In Section III, we introduce the notion of natural super-orbitals for operators and define orbital rotations in terms of super-orbitals in a doubled Hilbert space. We also analyze the properties of such rotations when the operator is unitary, and illustrate the formalism with a simple case of a factorized operator. In Section IV, we consider two limiting cases: non-interacting systems, where the natural orbital basis trivializes the MPO structure, and Haar-random unitaries, where no basis is preferred. The main numerical results are presented in Section V, where we investigate the time-evolution operator and time-evolved local observables in the $t$$V$ model and in a quantum impurity model. We show that in the latter case, the MPO complexity can be significantly reduced using the natural orbital basis. Finally, in Section VI, we summarize our findings and outline possible future directions stemming from the concepts introduced in this work.
@@ II One-body density matrix and natural orbitals for states
We start this article by reviewing the definition of the usual one-body density matrix defined for an arbitrary state $|\psi\rangle$, and some of its properties that we extend for operators later on. Given a fermionic system in which the particle can occupy $L$ orbitals of any kind (we only consider orbitals with occupation zero or one here), equipped with the fermionic creation and annihilation operators $\hat{c}^{\dagger}_{l}$ and $\hat{c}_{l}$ respectively creating and removing a fermion in the orbital $l$, a state $|\psi\rangle$ can be written in a complete basis of the ermionic Fock space:
@@
$Q_{ij}=\langle\psi|\hat{c}^{\dagger}_{i}c_{j}|\psi\rangle.$ (2)
@@
The property of the natural orbitals is that it is the basis in which the sum of the occupations $\sum_{a=1}^{r}n_{a}$ of the first $r$ orbitals sorted in descending order is maximized. This object is widely used in the quantum chemistry community to organize molecular orbitals according to their degree of correlations [36, 37, 38]. Therefore, if the state $|\psi\rangle$ can be represented as a product state in a given (known or unknown) one-particle basis, it will be revealed by the natural orbital basis, whose occupations will be strictly $n_{\alpha}=1$ or 00, if the orbital $\alpha$ is occupied by a fermion in the state ($n_{\alpha}=1$) or not ($n_{\alpha}=0$). Correlations lead to occupations that are different from zero or one.
@@ III Natural super-orbitals for operators
where the transpose is taken on the operators acting on the dual space labelled with a tilda, since they act from the left to the right in Eq. (5). The vacuum state is the tensor product of the vacua of the direct and dual spaces, $\rangle\!\rangle=|0\rangle\otimes|0\rangle$ | 0 ⟩, or equivalently the vacuum of a super-space corresponding to a doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. We can simplify the form of Eq. (6), by introducing the following super-operators acting on the doubled Hilbert space, which we label with a double hat:
@@
$\begin{split}R[\hat{O}]_{m,n}&=\langle\!\langle\hat{O}|\hat{\hat{d}}^{\dagger}_{m}\hat{\hat{d}}_{n}|\hat{O}\rangle\!\rangle\\ &=\langle\!\langle\hat{O}|\begin{pmatrix}\hat{c}^{\dagger}_{m}\hat{c}_{n}\otimes\mathds{1}&\hat{c}^{\dagger}_{m}\otimes\hat{\tilde{c}}_{n}\\ \hat{c}_{m}\otimes\hat{\tilde{c}}^{\dagger}_{n}&\mathds{1}\otimes\hat{\tilde{c}}_{m}\hat{\tilde{c}}^{\dagger}_{n}\end{pmatrix}|\hat{O}\rangle\!\rangle,\end{split}$ (9)
@@
$\begin{split}\langle\!\langle\hat{O}|\hat{c}^{\dagger}_{m}\hat{c}_{n}\otimes\mathds{1}|\hat{O}\rangle\!\rangle&=\mathrm{Tr}\left(\hat{O}^{\dagger}\hat{c}^{\dagger}_{m}\hat{c}_{n}\hat{O}\right),\\ \langle\!\langle\hat{O}|\hat{c}^{\dagger}_{m}\otimes\hat{\tilde{c}}_{n}|\hat{O}\rangle\!\rangle&=\mathrm{Tr}\left(\hat{O}^{\dagger}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n}\right),\\ \langle\!\langle\hat{O}|\mathds{1}\otimes\hat{\tilde{c}}_{m}\hat{\tilde{c}}^{\dagger}_{n}|\hat{O}\rangle\!\rangle&=\mathrm{Tr}\left(\hat{O}^{\dagger}\hat{O}\hat{c}_{m}\hat{c}^{\dagger}_{n}\right),\\ \end{split}$ (11)
@@
$\begin{split}\langle\!\langle\hat{O}|\hat{c}^{\dagger}_{m}\hat{c}_{n}\otimes\mathds{1}|\hat{O}\rangle\!\rangle&=\frac{\mathrm{Tr}(\mathds{1})}{2}\delta_{mn},\\ \langle\!\langle\hat{O}|\mathds{1}\otimes\hat{\tilde{c}}_{m}\hat{\tilde{c}}^{\dagger}_{n}|\hat{O}\rangle\!\rangle&=\frac{\mathrm{Tr}(\mathds{1})}{2}\delta_{mn},\\ \end{split},$ (12)
@@
where $A_{mn}=\mathrm{Tr}(\hat{O}^{\dagger}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n})$ is the only-non trivial $(L\times L)$ block of $R$. With such a form, the eigenvalues of $R$ can be found through the eigenvalues of $A^{\dagger}A$ only. If we denote the eigenvalues of $A^{\dagger}A$ as $\mu_{\alpha}$, the eigenvalues of $R$ come by pairs as $\lambda_{\alpha}^{\pm}=\frac{1}{2}\pm\sqrt{A^{\dagger}A}=\frac{1}{2}\pm|\mu_{\alpha}|$. If $A$ is normal (hermitian, unitary, etc. such that $A^{\dagger}A=AA^{\dagger}$), then $\mu_{\alpha}$ are the eigenvalues of $A$. The eigenvalues of $R$ being comprised between 00 and $1$, the eigenvalues of $A$ are thus restricted to the interval $[-\frac{1}{2},\frac{1}{2}]$. The eigenvectors of $R$ are just two copies of the eigenvectors of $A^{\dagger}A$ with a sign depending on the corresponding eigenvalue one for each off-diagonal block of the matrix: $W_{\alpha,\pm}=\frac{1}{\sqrt{2}}[v_{\alpha},\pm v_{\alpha}]$, where $v_{\alpha}$ is the eigenvector of $A^{\dagger}A$ associated to the eigenvalue $\mu_{\alpha}$. We recall that the doubled Hilbert space is divided into the operators acting on the direct space first, then on the dual space, as $\{\hat{c}^{\dagger}_{1},\hat{c}^{\dagger}_{2},\cdots,\hat{\tilde{c}}_{1},\hat{\tilde{c}}_{2},\cdots\}$. With the eigenvalues coming in pairs, the eigenvectors also come in pairs:
@@
$\rho=|\psi\rangle\langle\psi|\rightarrow|\rho\rangle\!\rangle=|\psi\rangle\otimes|\tilde{\psi}\rangle.$ (16)
with the normalization $\langle\!\langle\rho|\rho\rangle\!\rangle=\langle\psi|\psi\rangle\langle\tilde{\psi}|\tilde{\psi}\rangle=1$. Note that for a pure state, $|\tilde{\psi}\rangle=|\psi\rangle$, but it does not matter to keep them different. The corresponding matrix $R$ reads, if we write it by blocks involving the original fermion operators as follows:
@@
The terms not preserving the particle number like $\langle\psi|\hat{c}^{\dagger}_{n}|\psi\rangle$ are simply 00, such that:
$R[\rho]=\begin{pmatrix}Q[\tilde{\psi}]&0\\ 0&1-\tilde{Q}[\tilde{\psi}]\end{pmatrix}$ (18)
since $\tilde{Q}=Q$ for a pure state, the $R$ matrix is just twice the copy of the $Q$ matrix of the corresponding state when it’s particle-hole symmetric. For a mixed state holding a well defined charge, the matrices $Q$ and $\tilde{Q}$ are different but the dual and direct space do not mix as well. We can show the same results by using the definition of the elements of $R$ through the trace of operators as defined in Eq. (11) equivalently for density matrices. A pure state holds the following properties: $\rho^{\dagger}=\rho$ and $\rho^{\dagger}\rho=\rho$. In this case, only the diagonal blocks of the $R$ matrix are non-zero contrary to unitary operators and they read:
$\mathrm{Tr}\left[\rho\hat{c}^{\dagger}_{m}\hat{c}_{n}\right]=\langle\psi|\hat{c}^{\dagger}_{m}\hat{c}_{n}|\psi\rangle$ (19)
Correlation entropy
When the operator can be factorized as a product operators, the occupations of the natural super-orbitals are either 00 or $1$, and the correlation entropy will be $S_{\mathrm{corr}}=0$. In the opposite case, the furthest from a product operator an operator is, the more natural super-orbitals have their occupation $n_{\alpha}$ close to $\frac{1}{2}$, and the correlation entropy will be maximal when every orbital reaches this value, leading to the extensive maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$. We point out here that the spectrum of $R$ does not depend on the basis in which the operator is defined, operators which can be related through a one-body rotation (in the sense defined above) will have the same natural super-orbitals and thus the same correlation entropy.
@@
In order to illustrate the notion of product operator, we provide here a simple example of an operator that can be factorized and whose natural super-orbitals occupations are either 00 or $1$. In general, an operator can be written as a sum of projectors between different product states in the basis it is defined, as we wrote in Eq. (5). Usually, operators are defined in the computational basis $|i\rangle$, in which each index corresponds to a fermion in a site: $|101\rangle=\hat{c}^{\dagger}_{2}\hat{c}^{\dagger}_{0}|0\rangle$. We take here the example of the operator $\hat{c}^{\dagger}_{0}$, which we aim to write in a factorized form, for $L=3$:
@@
We accounted for the fermionic signs, and the effect of the operator on an incoming state is to add a fermion at site 00 when allowed. Its vectorized form looks like:
@@
[…]

### v2

@@ Abstract
We introduce the concept of natural super-orbitals for many-body operators, defined as the eigenvectors of the one-body super-density matrix associated with a vectorized operator. We relate these objects to measures of non-Gaussianity of operators associated to the occupations of the natural super-orbitals, and define how the non-stabilizerness of operators can be affected by such a basis rotation. We first analyze the general analytical properties of these objects in various contexts, including the time-evolution operator of non-interacting Hamiltonians and Haar-random unitaries. We then perform a numerical investigation of the natural super-orbitals corresponding to both the time-evolution operator and a time-evolved local operator, focusing on two many-body systems: the fermionic $t\text{-}V$ chain and an impurity model, using tensor network simulations. Our results reveal that the $t\text{-}V$ model lacks a preferred super-orbital basis, while in the impurity model, the occupations of the natural orbitals for both operators decay exponentially at all times. This indicates that only a small number of orbitals contribute significantly to quantum correlations, enabling a compact matrix-product-operator representation and a reduced non-stabilizerness in the natural orbital basis. Finally, we examine the spatial spread of the natural orbitals for time-evolved local operators in the impurity model and show that the complexity of this operator in the natural orbital basis saturates over time. This new framework opens the door to future research that leverages the compressed structure of operators in their natural super-orbital basis, enabling for instance the computation of out-of-time-order correlators in large interacting systems over extended time scales.
@@ I Introduction
Beyond its utility in quantifying and potentially reducing the complexity of classical simulations, the framework presented here also proves valuable for characterizing computational complexity in quantum simulations. A central topic of recent research has been the classification and quantification of quantum operations based on their resource requirements [36]. According to the Gottesman-Knill theorem, Clifford circuits can be simulated efficiently on classical computers, as they require only polynomial resources [37, 38]. However, these circuits cannot generate all quantum states. To achieve universal quantum computation, non-Clifford gates must be introduced, which significantly increases the difficulty of classical simulations [39, 40]. The complexity introduced by such non-Clifford operations is captured by the concept of non-stabilizerness [41, 42, 43, 44], also known as quantum magic, which has also can also be extended to fermionic systems [45, 46]. Quantifying this resource demands efficient and practical methods for classical computation [47, 48, 49, 50, 51]. Similarly, fermionic Gaussian states—eigenstates of non-interacting Hamiltonians or those generated by matchgate circuits—are also amenable to efficient classical simulation [52, 53]. In contrast, the inclusion of interactions introduces non-Gaussian features, thereby increasing complexity as the underlying computational Fock space is exponentially large. This other kind of complexity is captured by the measure of non-Gaussianity, which is interesting for different class of problems [54, 55, 56, 57, 46]. Both non-stabilizerness and non-Gaussianity become particularly meaningful in the context of operators, which represent quantum circuits or time-evolution operators associated to a given Hamiltonian [53, 58], and have been so far less investigated due to the lack of framework for operators. In this work, we demonstrate that these complexity measures can be extended to many-body operators and efficiently computed using tensor network methods. We also explore the properties of these measures in the specific systems studied here.
The paper is organized as follows: In Section II, we review the properties of the one-body density matrix for quantum states, focusing on the features we aim to generalize to operators. In Section III, we introduce the notion of natural super-orbitals for operators and define orbital rotations in terms of super-orbitals in a doubled Hilbert space. We also analyze the properties of such rotations when the operator is unitary, and illustrate the formalism with a simple case of a factorized operator. We then discuss the measures of non-Gaussianity and non-stabilizerness for operators. In Section IV, we consider two limiting cases: non-interacting systems, where the natural orbital basis trivializes the MPO structure, and Haar-random unitaries, where no basis is preferred. The main numerical results are presented in Section V, where we investigate the time-evolution operator and time-evolved local observables in the ttVV model and in a quantum impurity model. We show that in the latter case, the MPO complexity as well as the basis-dependent non-stabilizerness can be significantly reduced using the natural orbital basis. Finally, in Section VI, we summarize our findings and outline possible future directions stemming from the concepts introduced in this work.
@@ II One-body density matrix and natural orbitals for states
We start this article by reviewing the definition of the usual one-body density matrix defined for an arbitrary state $|\psi\rangle$, and some of its properties that we extend for operators later on. Given a fermionic system in which the particle can occupy $L$ orbitals of any kind (we only consider orbitals with occupation zero or one here), equipped with the fermionic creation and annihilation operators $\hat{c}^{\dagger}_{l}$ and $\hat{c}_{l}$ respectively creating and removing a fermion in the orbital $l$, a state $|\psi\rangle$ can be written in a complete basis of the fermionic Fock space:
@@
$Q_{ij}=\langle\psi|\hat{c}^{\dagger}_{i}\hat{c}_{j}|\psi\rangle.$ (2)
@@
The property of the natural orbitals is that it is the basis in which the sum of the occupations $\sum_{a=1}^{r}n_{a}$ of the first $r$ orbitals sorted in descending order is maximized. This object is widely used in the quantum chemistry community to organize molecular orbitals according to their degree of correlations [59, 60, 61]. Therefore, if the state $|\psi\rangle$ can be represented as a product state in a given (known or unknown) one-particle basis, it will be revealed by the natural orbital basis, whose occupations will be strictly $n_{\alpha}=1$ or 0, if the orbital $\alpha$ is occupied by a fermion in the state ($n_{\alpha}=1$) or not ($n_{\alpha}=0$). Correlations lead to occupations that are different from zero or one.
@@ III Natural super-orbitals for operators
where the transpose is taken on the operators acting on the dual space labelled with a tilda, since they act from the left to the right in Eq. (5). The vacuum state is the tensor product of the vacua of the direct and dual spaces, $|0\rangle\!\rangle=|0\rangle\otimes|0\rangle$, or equivalently the vacuum of a super-space corresponding to a doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. We can simplify the form of Eq. (6), by introducing the following super-operators acting on the doubled Hilbert space, which we label with a double hat:
@@
[…]
@@
[…]
@@
[…]
@@
where $A_{mn}=\mathrm{Tr}(\hat{O}^{\dagger}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n})$ is the only-non trivial $(L\times L)$ block of $R$. With such a form, the eigenvalues of $R$ can be found through the eigenvalues of $A^{\dagger}A$ only. If we denote the eigenvalues of $A^{\dagger}A$ as $\mu_{\alpha}$, the eigenvalues of $R$ come by pairs as $\lambda_{\alpha}^{\pm}=\frac{1}{2}\pm\sqrt{A^{\dagger}A}=\frac{1}{2}\pm|\mu_{\alpha}|$. If $A$ is normal (hermitian, unitary, etc. such that $A^{\dagger}A=AA^{\dagger}$), then $\mu_{\alpha}$ are the eigenvalues of $A$. The eigenvalues of $R$ being comprised between 0 and $1$, the eigenvalues of $A$ are thus restricted to the interval $\frac{1}{2},\frac{1}{2}]$. The eigenvectors of $R$ are just two copies of the eigenvectors of $A^{\dagger}A$ with a sign depending on the corresponding eigenvalue one for each off-diagonal block of the matrix: $W_{\alpha,\pm}=\frac{1}{\sqrt{2}}[v_{\alpha},\pm v_{\alpha}]$, where $v_{\alpha}$ is the eigenvector of $A^{\dagger}A$ associated to the eigenvalue $\mu_{\alpha}$. We recall that the doubled Hilbert space is divided into the operators acting on the direct space first, then on the dual space, as $\{\hat{c}^{\dagger}_{1},\hat{c}^{\dagger}_{2},\cdots,\hat{\tilde{c}}_{1},\hat{\tilde{c}}_{2},\cdots\}$. With the eigenvalues coming in pairs, the eigenvectors also come in pairs:
@@
$\hat{\rho}=|\psi\rangle\langle\psi|\rightarrow|\hat{\rho}\rangle\!\rangle=|\psi\rangle\otimes|\tilde{\psi}\rangle$. (16)
with the normalization $\langle\!\langle\hat{\rho}|\hat{\rho}\rangle\!\rangle=\langle\psi|\psi\rangle\langle\tilde{\psi}|\tilde{\psi}\rangle=1$. Note that for a pure state, $|\tilde{\psi}\rangle=|\psi\rangle$, but it does not matter to keep them different. The corresponding matrix $R$ reads, if we write it by blocks involving the original fermion operators as follows:
@@
The terms not preserving the particle number like $\langle\psi|\hat{c}^{\dagger}_{n}|\psi\rangle$ are simply 0, such that:
[…]
since $\tilde{Q}=Q$ for a pure state, the $R$ matrix is just twice the copy of the $Q$ matrix of the corresponding state when it’s particle-hole symmetric. For a mixed state holding a well defined charge, the matrices $Q$ and $\tilde{Q}$ are different but the dual and direct space do not mix as well. We can show the same results by using the definition of the elements of $R$ through the trace of operators as defined in Eq. (11) equivalently for density matrices. A pure state holds the following properties: $\hat{\rho}^{\dagger}=\hat{\rho}$ and $\hat{\rho}^{\dagger}\hat{\rho}=\hat{\rho}$. In this case, only the diagonal blocks of the $R$ matrix are non-zero contrary to unitary operators and they read:
$\mathrm{Tr}\left[\hat{\rho}\hat{c}^{\dagger}_{m}\hat{c}_{n}\right]=\langle\psi|\hat{c}^{\dagger}_{m}\hat{c}_{n}|\psi\rangle$ (19)
@@
In order to illustrate the notion of product operator, we provide here a simple example of an operator that can be factorized and whose natural super-orbitals occupations are either 0 or $1$. In general, an operator can be written as a sum of projectors between different product states in the basis it is defined, as we wrote in Eq. (5). Usually, operators are defined in the computational basis $|i\rangle$, in which each index corresponds to a fermion in a site: $|101\rangle=\hat{c}^{\dagger}_{2}\hat{c}^{\dagger}_{0}|0\rangle$. We take here the example of the operator $\hat{c}^{\dagger}_{0}$, which we aim to write in a factorized form, for $L=3$:
@@
We accounted for the fermionic signs, and the effect of the operator on an incoming state is to add a fermion at site 0 when allowed. Its vectorized form looks like:
@@
$\begin{split}\langle\!\langle\hat{\hat{n}}_{1}\rangle\!\rangle&=1,\langle\!\langle\hat{\hat{n}}_{3}\rangle\!\rangle=0,\langle\!\langle\hat{\hat{n}}_{5}\rangle\!\rangle=0,\\ 1-\langle\!\langle\hat{\hat{n}}_{2}\rangle\!\rangle&=1,\;1-\langle\!\langle\hat{\hat{n}}_{4}\rangle\!\rangle=1,\;1-\langle\!\langle\hat{\hat{n}}_{6}\rangle\!\rangle=1\\ &\equiv[0,0,1,1,1,1],\end{split}$ (24)
@@
Correlation entropy
When the operator can be factorized as a product operators, the occupations of the natural super-orbitals are either 0 or $1$, and the correlation entropy will be $S_{\mathrm{corr}}=0$. In the opposite case, the furthest from a product operator an operator is, the more natural super-orbitals have their occupation $n_{\alpha}$ close to $\frac{1}{2}$, and the correlation entropy will be maximal when every orbital reaches this value, leading to the extensive maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$. We point out here that the spectrum of $R$ does not depend on the basis in which the operator is defined, operators which can be related through a one-body rotation (in the sense defined above) will have the same natural super-orbitals and thus the same correlation entropy.
Operator non-Gaussianity and non-stabilizerness
[…]
Non-stabilizerness: Characterizing the non-stabilizerness of an operator provides insight into its computational complexity, particularly in terms of tableau classical simulation [64, 38] and of the number of non-Clifford gates required for its realization on a quantum device. A standard measure of this resource is the Operator Stabilizer Rényi Entropy (OSE) [47, 46], defined as:
$\begin{split}\mathcal{M}^{(\alpha)}(\hat{O})&=\frac{1}{1-\alpha}\log\left(P^{(\alpha)}(\hat{O})\right)\\ P^{(\alpha)}(\hat{O})&=\sum\limits_{\mathcal{S}\in\mathcal{S}_{L}}\!\!\left(\frac{|\mathrm{Tr}\left(\hat{O}^{\dagger}\mathcal{S}\right)\!\!|}{2^{L}}\right)^{\!\!2\alpha}\!\!\!\!=\sum\limits_{\mathcal{S}\in\mathcal{S}_{L}}|a_{\mathcal{S}}|^{2\alpha},\end{split}$ (26)
where the coefficients $a_{\mathcal{S}}$ arise from the expansion $\hat{O}=\sum\limits_{\mathcal{S}\in\mathcal{S}_{N}}a_{\mathcal{S}}\hat{\mathcal{S}}$, which can be defined for the fermionic degrees of freedom as stated in the previous paragraph, $\mathcal{S}_{L}=\{\mathds{1},\hat{c}_{i}^{\dagger},\hat{c}_{i},\hat{n}_{i}\}^{\otimes L}$. Usually, $\alpha=2$ is considered. It measures the number of significant product operators (or Pauli strings when working in the computational basis) on which the operator is decomposed. In contrast to the non-Gaussianity, the non-stabilizerness is basis-dependent. For a Gaussian operator, we previously showed that it decomposes into a single product operator in the natural super-orbital basis, yielding vanishing non-stabilizerness. This basis also maximizes the weight of the coefficients $a_{\mathcal{S}}$ [59] given a number $r\leq 2L$ of super-orbitals included, thereby minimizing the non-stabilizerness under Gaussian transformations. Notably, when certain orbitals factorize in this basis, they do not contribute to the sum in Eq. (26), resulting in an exponential reduction in the number of terms and thus in the non-stabilizerness. Such behavior has been observed in the evolution of local operators that exhibit a light-cone structure, where orbitals outside the light-cone effectively decouple [58]. We will demonstrate that similar behavior arises in the time-evolution operator of impurity models when expressed in the natural super-orbital basis.